This is the exact question asked:
Give an equation of the parabola which opens either up or down given the three points of: (0,11), (1,21), and (-2,3) on the parabola.
I just don't know how to do this! Thanks so much and God bless you.
-Zach
Answer by stanbon(75887) 
 Give an equation of the parabola which opens either up or down given the three points of: (0,11), (1,21), and (-2,3) on the parabola.
---------------
Form: ax^2 + bx + c = y
-----
Substitute each point into the Form and solve for a,b,c.
-------------------------------
0 + 0 + c = 11
a + b + c = 21
4a-2b + c = 3
----------------------
Use any method you know to get:
a = 2
b = 8
c = 11
----------
Equation:
y = 2x^2 + 8x + 11
==========================
Cheers,
Stan H.
| |
|